20cm3 of a gaseous Hydrocarbon were mixed with 90cm3 of O2, and the mixture exploded. At room tempt 60cm3 of gas were left. 40cm3 of the gas(CO2) were adsorbed by NaOH, leaving 20cm3 of O2. Find the empirical formula of the hydrocarbon
1
Expert's answer
2019-09-09T05:00:59-0400

"C_xH_y + (x+y/4) O_2 → x CO_2 +\frac{y}{2} H_2O"

As per Avagadro law,same volume of gases consist of same number of moles.

So,"x(vol \space of \space CO_2)=40 cm^3"

Volume of oxygen used"=90-20=70 cm^3"

As per stoichiometry,

"x+\frac{y}{4}=70"

"40+\frac{y}{4}=70 \implies y=120"

Emperical formula"=C_xH_y=C_{70}H_{120}=C_7H_{12}"